### v1

@@ Abstract
The introduction of blobs through EIP-4844 has significantly reduced the Data Availability (DA) costs for rollups on Ethereum. However, due to the fixed size of blobs at 128 KB, rollups with low data throughput face a dilemma: they either use blobs inefficiently or decrease the frequency of DA submissions. Blob sharing, where multiple rollups share a single blob, has been proposed as a solution to this problem. This paper examines the effectiveness of blob sharing based on real-world data collected approximately six months after the implementation of EIP-4844. By simulating cost changes using a simple blob sharing format, we demonstrate that blob sharing can substantially improve the costs and DA service quality for small rollups, effectively resolving their dilemma. Notably, we observed cost reductions in USD exceeding 90% for most of the rollups when they cooperate, attributable to the smoothing effect of the blob base fee achieved through blob sharing.
@@ 2.2 Simulation Model
    Handling Excess Data: If rollup data exceeds 128 KB, the surplus data is configured as a new blob.
@@
    Blob Submission Limit per Block: If the number of blobs generated in the current block exceeded six, the excess blobs were deferred and submitted in the next block. If six or fewer blobs were generated, they were bundled into a transaction and submitted within the current block.
  5. 5.

### v2

@@ Abstract
The introduction of blobs through EIP-4844 has significantly reduced the Data Availability (DA) costs for rollups on Ethereum. However, due to the fixed size of blobs at 128 KB, rollups with low data throughput face a dilemma: they either use blobs inefficiently or decrease the frequency of DA submissions. Blob sharing, where multiple rollups share a single blob, has been proposed as a solution to this problem. This paper examines the effectiveness of blob sharing based on real-world data collected approximately six months after the implementation of EIP-4844. By simulating cost changes using a simple blob sharing format, we demonstrate that blob sharing can substantially improve the costs and DA service quality for small rollups, effectively resolving their dilemma. Notably, we observed cost reductions in USD exceeding 85% for most of the rollups when they cooperate, attributable to the smoothing effect of the blob base fee achieved through blob sharing.
@@ 2.2 Simulation Model
    Handling Excess Data/Blob: If rollup data exceeds 128 KB, the surplus data is configured as a new blob. And if the number of blobs generated in the current block exceeded six, the excess blobs were deferred and submitted in the next block. If six or fewer blobs were generated, they were bundled into a transaction and submitted within the current block.